Determine Cash Flows

Natural Foods Inc. is planning to invest in new manufacturing equipment to make a new garden tool. The new garden tool is expected to generate additional annual sales of 5,300 units at $30 each. The new manufacturing equipment will cost $63,100 and is expected to have a 10-year life and a $4,800 residual value. Selling expenses related to the new product are expected to be 4% of sales revenue. The cost to manufacture the product includes the following on a per-unit basis:

Direct labor $5.10
Direct materials 16.70
Fixed factory overhead-depreciation 1.10
Variable factory overhead 2.60
Total $25.50

Determine the net cash flows for the first year of the project, Years 2–9, and for the last year of the project. Use the minus sign to indicate cash outflows. Do not round your intermediate calculations but, if required, round your final answers to the nearest dollar.

Natural Foods Inc.
Net Cash Flows
Year 1 Years 2-9 Last Year
Initial investment $
Operating cash flows:
Annual revenues $ $ $
Selling expenses
Cost to manufacture
Net operating cash flows $ $ $
Total for Year 1 $
Total for Years 2–9 (operating cash flow) $
Residual value
Total for last year $
Answer #1
Natural Foods Inc.
Net Cash Flows
Year 1 Years 2-9 Last Year
Initial investment -63100
Operating cash flows:
Annual revenues 159000 159000 159000
Selling expenses -6360 -6360 -6360
Cost to manufacture -129320 -129320 -129320
Net operating cash flows 23320 23320 23320
Total for Year 1 -39780
Total for Years 2–9 (operating cash flow) 186560
Residual value 4800
Total for last year 28120
Workings:
Annual revenues 159000 =5300*30
Selling expenses 6360 =159000*4%
Cost to manufacture 129320 =5300*(25.50-1.10)